OpenStudy (anonymous):

Use the alternating series estimation theorm to determine how many terms should be used to estimate the sum of the entire series with an error of less than 0.001.

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} (-1)^{n+1}\frac{ 1 }{ n+\sqrt{5n}}^{3}\]

OpenStudy (anonymous):

@amistre64 :P

OpenStudy (amistre64):

if we need to use a thrm, define the thrm

OpenStudy (anonymous):

I only know the alternating series thrm, I dont understand the estimation one.

OpenStudy (amistre64):

ive never heard of the esimating one so the definition will be required to see if i can even understand it :)

OpenStudy (anonymous):

Lol okay, I'll figure it out. Thanks!

OpenStudy (amistre64):

that link doesnt show for me, guess ill find it anothe rplace tho

OpenStudy (amistre64):

https://www.youtube.com/watch?v=ilRkVcAxdCo here he is on the youtube ...

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

looks to me that we simply see when the 'rule' is equal to .001 and solve for n

OpenStudy (anonymous):

The answer i got is 0.0295

OpenStudy (anonymous):

which is the first term..

OpenStudy (amistre64):

\[ \left|\frac{ 1 }{ n+\sqrt{5n}}^{3}\right|<=.001\] \[ \frac{ 1 }{ n+\sqrt{5n}}<=\sqrt[3]{.001}\] \[ \frac{ 1 }{\sqrt[3]{.001} }<=n+\sqrt{5n}\] ... maybe

OpenStudy (amistre64):

.0295 is not < .001

OpenStudy (amistre64):

run some trials, if n=10 < .001, try n=5 if n=5 > .001, try n=7 and so on to narrow in on it

OpenStudy (anonymous):

Well what patrick did, was he calculated the first couple terms, and then whatever number was = to decimal places, he added the numbers before it.

OpenStudy (anonymous):

exlcuding the number that fit the decimal place.

OpenStudy (amistre64):

right, i watched the first example :) 10 minutes is too long lol

OpenStudy (amistre64):

but the estimation thrm says that the error is at most the size of the first neglected term

OpenStudy (amistre64):

i get n=5

OpenStudy (amistre64):

or 4 depending on how you look at it the first term less then .001 i when n=5

OpenStudy (amistre64):

yeah brute math instead of solving algebraically

OpenStudy (anonymous):

Alright, well thanks :P

OpenStudy (amistre64):

the 5th term is equal to .001 :) so the first 4 terms shold have an error that is no more than .001

OpenStudy (anonymous):

Okay, thanks so much!

OpenStudy (amistre64):

youre welcome